how do i write y=2x^2+4x-3 in the form of y=a(x-h)^2+k
1
Expert's answer
2014-07-07T04:36:52-0400
y=2x2+4x-3=2(x2+2x)-3=2(x2+2x+1-1)-3=2(x2+2x+1)-2-3=2(x+1)2-5